OpenStudy (anonymous):

12t^2+17t=40 -40 -40 -40+12t^2+17t (-8-3t) (5-4t) -40+32t-15+12t^2 -40+17t+12t^2 i should have set each factor to zero..help

OpenStudy (anonymous):

I don't know the question, so I will just guess the question is to find the value of t. \[12t^2+17t=40\]\[12t^2+17t-40=0\]\[(3t+8)(4t-5)=0\] \[3t+8=0\]or \[4t-5=0\] \[3t=-8\]or \[4t=5\] \[t=-8/3\]or \[t=5/4\]